# The Too Visible QCD Axion Luca Di Luzioa, Michele Redib, Alessandro Strumiac, Andrea Tesib, Arsenii V. Titovd,a

### v1

@@ 1 Introduction
In a recent paper Murayama proposed a novel axion theory to address the strong CP problem [1], where all the relevant dynamics takes place at the QCD scale. As well known, QCD solves the strong CP problem if the bare mass of the up quark vanishes, $m_{u}=0$. In this case the $\eta^{\prime}$ acts as an axion, dynamically relaxing the $\theta$-angle to zero. However, this possibility is excluded by comparing lattice computations with data [2]. Building on the massless up quark solution, Murayama added a new complex scalar Standard Model (SM) singlet $\varphi$ coupled to up quarks with unit Peccei-Quinn (PQ) charge under the rotation of $u_{R}$. When quarks develop their condensate, $\varphi$ acquires a PQ-breaking vacuum expectation value effectively giving a mass to the up quark. The main issue is that the $\pi^{0}$ mass is modified
@@
The observed pion mass splitting $m_{\pi^{\pm}}-m_{\pi^{0}}\approx 0.03m_{\pi^{0}}$ arises predominantly from electromagnetic self-energy effects, with only a small correction from the up-down quark mass difference. Lattice computations claim a per-cent level precision on the electromagnetic effect [3, 4, 5]. So this model is excluded unless the correction to the neutral pion mass can be nearly totally canceled.
We try improving on this situation. In section 2 we review and extend the model. In section 3 we find some improvement, but only partial. Neither the Kaplan-Manohar ambiguity nor one-loop corrections to pion masses avoid the problem. The pion mass splitting is only mildly reduced by assuming that the down quark $d_{R}$ too is PQ charged. In view of $m_{d}/m_{u}\approx 2$, the pion mass splitting could be avoided by assuming PQ charges $n_{u}/n_{d}=2$; however the new physics gets non-perturbative below the QCD scale. This can be cured by adding new light PQ scalars, but scans in the larger parameter space do not find an allowed solution. Isospin is recovered in a model where the $s$ quark only is PQ charged, but the new PQ scalars turn out to be too light.
In section 4 we identify a structural obstacle. Like quark masses, the interactions of $\varphi$ break the accidental SU(3)u,d,s flavour symmetry of massless QCD. As a result, integrating out QCD and $\varphi$, gives an effective chiral Lagrangian for mesons $U$ that differs from the usual QCD chiral Lagrangian. The operator related to quark masses is substituted by a different higher-order operator that needs to have comparable size,
@@
containing the new PQ spurion $I_{\rm PQ}$, given by $I_{\rm PQ}\propto\mathrm{diag}(m_{u}\,,0\,,0)$ in the model of [1]. This spurion breaks the SU(2)u,d isospin symmetry in a higher-order chiral operator, so pion masses get the order one correction of eq. (1), and pion scatterings too are similarly affected. The general problem of the new theories is that they unavoidably predict order unity corrections to meson physics, but the QCD chiral Lagrangian has been more precisely tested.
@@ 2 The model and its generalizations
We start our discussion reviewing and extending the results in [1].
@@ 2.1 General PQ charges
We generalise the theory allowing for general PQ charges $n_{q}$ and couplings of $\varphi$ to light quarks $q$, including down and strange quarks. The effective Lagrangian before QCD confinement reads
$\begin{split}\mathscr{L}&=-\frac{1}{4}G^{a}_{\mu\nu}G^{a,\mu\nu}+\theta\frac{g_{s}^{2}}{32\pi^{2}}G^{a}_{\mu\nu}\tilde{G}^{a,\mu\nu}+f_{\varphi}^{2}\bigg[|\partial_{\mu}\varphi|^{2}-m^{2}_{\varphi}|\varphi|^{2}-\lambda f_{\varphi}^{2}|\varphi|^{4}\bigg]+\\ &-\sum_{q=\{u,d,s\}}\bigg(m_{q}\delta_{n_{q}0}+\frac{f_{\varphi}\kappa_{q}}{\sqrt{2}}(\sqrt{2}\varphi)^{n_{q}}\bigg)\bar{q}_{L}q_{R}+\text{h.c.}\end{split}$ (3)
@@
PQ gets spontaneously broken by the QCD chiral condensate after confinement. On the minimum, the quarks charged under PQ acquire non-vanishing masses
$m_{q}=\kappa_{q}\frac{f_{\varphi}}{\sqrt{2}}\,.$ (5)
The quark couplings can then be rewritten as
$\sum_{q=\{u,d,s\}}m_{q}(\sqrt{2}\varphi)^{n_{q}}\bar{q}_{L}q_{R}+\hbox{h.c.}$ (6)
for both $n_{q}=0$ and $n_{q}\neq 0$.
@@
$\begin{split}\mathscr{L}_{\pi\varphi}&=f_{\varphi}^{2}\bigg[|\partial_{\mu}\varphi|^{2}-m^{2}_{\varphi}|\varphi|^{2}-\lambda f_{\varphi}^{2}|\varphi|^{4}\bigg]\\ &+\frac{f_{\pi}^{2}}{4}\operatorname{tr}[D_{\mu}UD^{\mu}U^{\dagger}]+\frac{f_{\pi}^{2}B_{0}}{2}\operatorname{tr}[M_{q}U^{\dagger}+M_{q}^{\dagger}U]-\frac{m_{1}^{2}f_{\pi}^{2}}{12}\left[\left(-\frac{i}{2}\ln\det U+\hbox{h.c.}\right)-\theta\right]^{2}.\end{split}$ (7)
@@
$\sigma$ is a radial mode, and $a$ is the angular mode related to the axion. The PQ symmetry allows to remove the $\theta$ angle and take the vacuum expectation values of the neutral mesons to be zero. The minimum condition comes from requiring $\sigma=0$ and $a=0$ on the vacuum. The value of $m_{\varphi}$ needed to get $\sigma=0$ is
@@
Since we have assumed that both $m_{\varphi}^{2}$ and $\lambda$ are positive, [^2] If the squared mass term is negative, PQ breaks spontaneously irrespectively of the QCD dynamics, and the model reduces to a DFSZ-like axion model [8, 9]. this model has an upper bound on $f_{\varphi}$. For example, focussing on the model $\{n_{u},n_{d},n_{s}\}=\{1,0,0\}$, we get
@@
A large decay constant can only be achieved for a tiny quartic. If $f_{\varphi}$ is well within this upper bound, the minimization condition is independent from $\lambda$, and $m_{\varphi}^{2}$ just depends on the mass of the heaviest quark with PQ charge. In the case of [1], we have [^3] Note that $m_{\varphi}$ is $\sqrt{2}$ times smaller than in [1]. This originates from the normalization of kinetic terms in the chiral Lagrangian.
@@
The mass of the radial mode is
@@
We next compute the masses of the CP-odd scalars, $a$ and the mesons $\pi^{0},\eta,\eta^{\prime}$.
@@ 2.2 CP-odd scalar mass matrix
Neglecting interactions with the CP-even $\sigma$, which amounts to set $\sigma=0$, the Lagrangian restricted to self-interactions of the CP-odd scalars around the minimum of (11) is
@@
It is useful to see explicitly how the strong CP problem is solved in this setup. Under a PQ transformation $\bar{q}_{L}q_{R}\to e^{i\alpha Q_{\rm PQ}}\bar{q}_{L}q_{R}$, with $Q_{\rm PQ}=\text{diag}(n_{u},n_{d},n_{s})$, the pseudo-scalar fields transform as: $a\to a-\alpha$ and $U\to Ue^{-i\alpha Q_{\rm PQ}}$. In particular
@@
The mixing of $a$ with the other states depends on the charges $n_{u,d,s}$ and on $f_{\pi}/f_{\varphi}$. If all quark charges are set to zero $a$ is massless. For large $f_{\varphi}\gg{\rm GeV}$, $a$ becomes weakly coupled to QCD as in the usual axion theories. In general, $f_{\varphi}$ is given by eq. (11). For charges $n_{q}\leq 1$, the (1,1) entry of the matrix can be rewritten as
@@ 2.3 Kaplan-Manohar ambiguity of the chiral Lagrangian
in the LO operator $\operatorname{tr}(M_{q}U^{\dagger})+\hbox{h.c.}$ [10], in view of $2\det M_{q}\,\operatorname{tr}M_{q}^{-1}=(\operatorname{tr}M_{q})^{2}-\operatorname{tr}M_{q}^{2}$. The shift term has the same symmetry transformations as $M_{q}$. The presence of $\det U=e^{i\sqrt{6}\eta^{\prime}/f_{\pi}}$ guarantees that the two terms have the same transformation also under the axial U(1)A. The constant $x_{\rm KM}$ is approximately determined by lattice computations. This is known as the Kaplan-Manohar (KM) ambiguity [10], since it complicates the extraction of bare quark masses from meson data. In particular, it replaces $m_{u}\to m_{u}+x_{\rm KM}m_{d}m_{s}$. This was considered as a loophole for reviving the massless up-quark solution to the strong CP problem, before lattice calculations established that $m_{u}\neq 0$ [11]. In our case the KM ‘ambiguity’ effectively redefines the quark mass matrix into
@@ 3 QCD-scale axion
The chiral Lagrangian at LO predicts that pions form a degenerate triplet (up to a tiny quark-mass-induced splitting), with the observed mass difference arising primarily from electromagnetic effects [3, 4, 5]
@@
where $\Lambda\sim m_{\rho}$. Adding a QCD-scale axion generically breaks the isospin symmetry, ruining this successful prediction.
@@ 3.1 Tree-level corrections to meson masses
This class of models predicts important modifications to meson masses. To simplify formulas, we now focus on $n_{s}=0$. In the limit where the $a$ is heaviest state, i.e. $m_{\varphi}>$ GeV, we can integrate it out. This adds to the QCD squared mass matrix of neutral mesons the following additive correction
@@ 3.2 Model with axion coupled to the up quark
Murayama [1] considered the case where the up quark only is PQ-charged, $n_{u}=1$ and $n_{d}=0$. In this case one finds (neglecting electro-magnetic corrections)
@@
that is equivalent to eq. (1). A QCD-scale $\varphi$ needs a relatively large Yukawa coupling, $\kappa_{u}\gtrsim 0.5$, not much below the perturbativity limit. The Kaplan-Manohar ambiguity helps to alleviate the problem with pion masses [1], but only mildly. It shifts the quark matrix into
@@
The KM shift terms do not modify $\langle{\varphi}\rangle$. The problematic correction to the pion mass gets suppressed if the up entry of $M_{q}|_{\rm KM}$ is dominated by the KM shift, as it is $\varphi$-independent. Assuming that the up quark mass induced by $\langle\varphi\rangle$ is a factor $R$ smaller than the measured up quark mass, we obtain
@@
Avoiding the pion mass problem requires $R\gtrsim 10$. This corresponds to a value of $x_{\rm KM}$ much larger than the value suggested by lattice computations, that favor a mild KM shift and disfavor reaching $m_{u}=0$, corresponding to $R=\infty$ [12, 13, 2]. Thus it seems unlikely that the KM ambiguity solves the pion mass problem.
If the model were viable, the new GeV-scale scalars should be identified with some observed hadron. Table 1 lists some candidates. Decay rates of $a$ and $\sigma$ can be estimated either from decays to quarks or to pions. Both computations are approximate and lead to similar estimates:
@@
The relatively narrow $a$ could be identified with the observed $\eta(1295)$, as its mass is $m_{a}\sim 1.1\,{\rm GeV}$ and its width is $\Gamma_{a}\sim 25\,{\rm MeV}$ for $\kappa_{u}\sim 1/2$ (we fixed factors of 2 compared to [1]). The broad $\sigma$ can be more easily hidden inside QCD hadron resonances.
@@ 3.3 Model with axion coupled to up and down quarks
Given that the pion mass splitting originates from the breaking of the SU(2) isospin symmetry, it is tempting to assume that both up and down quarks are PQ charged. For negligibile $\lambda$ the mass of $\sigma$ is
@@
A cancellation of the pion mass shift can happen for $n_{u}=2$ and $n_{d}=1$ (see eq. (25)), as data allow for $m_{d}/m_{u}=2$ within uncertainties. However $n_{u}=2$ means that the up-quark mass comes from a dimension-5 operator that gets strongly coupled at energies mildly above $m_{u}$, below the QCD scale. This would require a UV completion at that scale, with extra particles that would be problematically light.

### v2

@@ 1 Introduction
In a recent paper Murayama proposed a novel axion theory to address the strong CP problem [1], where all the relevant dynamics takes place at the QCD scale. As well known, QCD solves the strong CP problem if the bare mass of the up quark vanishes, $m_{u}=0$. In this case the $\eta^{\prime}$ acts as an axion, dynamically relaxing the $\theta$-angle to zero. However, this possibility is excluded by comparing lattice computations with data [2]. Building on the massless up quark solution, Murayama added a new complex scalar Standard Model (SM) singlet $\varphi$ coupled to up quarks with unit Peccei-Quinn (PQ) charge under the rotation of $u_{R}$. When quarks develop their condensate, $\varphi$ acquires a PQ-breaking vacuum expectation value effectively giving a mass to the up quark. The main issue of the proposal is that the $\pi^{0}$ mass is modified
@@
The observed pion mass splitting $m_{\pi^{\pm}}-m_{\pi^{0}}\approx 0.03m_{\pi^{0}}$ arises predominantly from electromagnetic self-energy effects, with only a small correction from the up-down quark mass difference. Since lattice computations claim a per-cent level precision on the electromagnetic effect [3, 4, 5], this model is excluded unless the correction to the neutral pion mass can be almost completely canceled.
In this work we try to improve this conclusion. In section 2 we review and extend the model to general PQ charges. In section 3 we discuss models that could improve the mass splitting, but only with partial success. Neither the Kaplan-Manohar ambiguity nor one-loop corrections to pion masses seem to avoid the problem. The pion mass splitting is only mildly reduced by assuming that the down quark $d_{R}$ has also PQ charge. In view of $m_{d}/m_{u}\approx 2$, the pion mass splitting could be avoided by assuming PQ charges $n_{u}/n_{d}=2$; however the new physics gets non-perturbative below the QCD scale in that case. This could be cured by adding new light PQ scalars, but scans in the larger parameter space do not find an allowed solution. Isospin invariance is recovered in a model where the $s$ quark only is PQ charged, but the new PQ scalars turn out to be too light.
In section 4 we identify a structural obstacle to build viable models. Like quark masses, the couplings of $\varphi$ break the accidental SU(3)u,d,s flavour symmetry of massless QCD. As a result, integrating out QCD and $\varphi$, gives an effective chiral Lagrangian for mesons $U$ that differs from the usual QCD chiral Lagrangian. The operator related to quark masses is replaced by a different higher-order operator of comparable size,
@@
containing the new PQ spurion $I_{\rm PQ}$, given by $I_{\rm PQ}\propto\mathrm{diag}(m_{u}\,,0\,,0)$ in the model of [1]. This spurion breaks the accidental SU(2)u,d isospin symmetry of the leading order chiral Lagrangian, so pion masses get the order one correction of eq. (1), and pion scatterings too are similarly affected. The general problem of the new theories is that they unavoidably predict order one corrections to meson physics, but the QCD chiral Lagrangian has been more precisely tested.
@@ 2 The model and its generalizations
We begin our discussion reviewing and extending the results in [1].
@@ 2.1 General PQ charges
We extend the SM introducing a complex scalar field with unit charge under a global U(1)PQ symmetry under which the light quarks $q\equiv(u\,,d\,,s)$ have charge $n_{q}\equiv(n_{u}\,,n_{d}\,,n_{s})$. The effective Lagrangian invariant under PQ symmetry (up to anomaly) at energies above QCD confinement and below electro-weak symmetry breaking reads
@@
For later convenience we introduced the dimension-less $\varphi$ complex field with decay constant $f_{\varphi}$. Field re-phasings allow to make $m_{q}$ real and positive. If one $n_{q}=0$, the corresponding quark $q$ acquires the usual mass term $m_{q}$. If instead $n_{q}\neq 0$ it acquires a coupling to $\varphi$. The PQ symmetry gets spontaneously broken by the QCD chiral condensate after confinement: quarks develop a condensate $\langle\bar{q}_{Li}q_{Rj}\rangle=-\delta_{ij}B_{0}f_{\pi}^{2}/2$, triggering a vacuum expectation value for $\varphi$. The mass parameter $m_{\varphi}^{2}$ will be fixed such that $\langle\varphi\rangle\equiv 1/\sqrt{2}$ in the vacuum. As a consequence, quarks acquire an effective mass $m_{q}$ from their coupling to $\varphi$. We write it
$m_{q}=\kappa_{q}\frac{f_{\varphi}}{\sqrt{2}}\,.$ (4)
The model [1] corresponds to $n_{u}=1$ and $n_{d}=n_{s}=0$. In this case $\kappa_{u}$ is a Yukawa coupling and the quark condensate induces a PQ-breaking tadpole. We consider a general PQ charge assignment of quarks. PQ charges $|n_{q}|>1$ correspond to a non-renormalizable interaction (with $\varphi\to\varphi^{*}$ if $n_{q}<0$). The case $n_{q}=2$ produces a PQ-breaking negative contribution to the squared mass of the real/radial part of $\varphi$.
@@
$\begin{split}\mathscr{L}_{\pi\varphi}&=f_{\varphi}^{2}\bigg[|\partial_{\mu}\varphi|^{2}-m^{2}_{\varphi}|\varphi|^{2}-\lambda f_{\varphi}^{2}|\varphi|^{4}\bigg]\\ &+\frac{f_{\pi}^{2}}{4}\operatorname{tr}[D_{\mu}UD^{\mu}U^{\dagger}]+\frac{f_{\pi}^{2}B_{0}}{2}\operatorname{tr}[M_{q}U^{\dagger}+M_{q}^{\dagger}U]-\frac{m_{1}^{2}f_{\pi}^{2}}{12}\left[\left(-\frac{i}{2}\ln\det U+\hbox{h.c.}\right)-\theta\right]^{2},\end{split}$ (6)
where the last term reproduces the QCD anomaly to leading order in $1/N$. Through a PQ transformation $\theta$ can be set to zero explaining why the strong interactions are CP invariant in the low energy effective theory.
@@
The radial mode $\sigma$ is a scalar while the phase $a$ is a pseudo-scalar. For $\theta=0$ (that can always be chosen through a PQ transformation) the vacuum expectation values of CP-odd scalars vanish due to CP invariance. $m_{\varphi}$ (or alternatively $f_{\varphi}$) is determined by the condition $\sigma=0$ in the minimum so that,
@@
Under the assumption that both $m_{\varphi}^{2}$ and $\lambda$ are positive, [^2] If the squared mass term is negative, PQ symmetry will be spontaneously broken independently of the QCD dynamics, and the model reduces to a DFSZ-like axion model [8, 9]. there is an upper bound on $f_{\varphi}$. For example, focusing on the model $\{n_{u},n_{d},n_{s}\}=\{1,0,0\}$, we get
@@
A large decay constant can only be achieved for a tiny quartic. If $f_{\varphi}$ is well within this upper bound, the minimization condition is independent of $\lambda$, and $m_{\varphi}^{2}$ is determined by the mass of the heaviest quark with PQ charge. In the case of [1], we have [^3] Note that $m_{\varphi}$ is $\sqrt{2}$ times smaller than in [1]. This discrepancy originates from the normalization of kinetic terms in the chiral Lagrangian.
@@
The mass of the radial mode is easily found
@@ 2.2 CP-odd scalar mass matrix
The mass of the pseudo-scalar $a$ is a little more complicated due to mixing with QCD mesons $\pi^{0},\eta,\eta^{\prime}$. Neglecting interactions with the CP-even $\sigma$, which amounts to set $\sigma=0$, the Lagrangian restricted to self-interactions of the CP-odd scalars around the minimum of (10) is
@@
It is useful to see explicitly how the strong CP problem is solved. Under a PQ transformation $\bar{q}_{L}q_{R}\to\bar{q}_{L}q_{R}e^{i\alpha Q_{\rm PQ}}$, with $Q_{\rm PQ}=\text{diag}(n_{u},n_{d},n_{s})$, the pseudo-scalar fields transform as: $a\to a-\alpha$ and $U\to e^{-i\alpha Q_{\rm PQ}}U$. In particular
@@
The mixing of $a$ with the other states depends on the charges $n_{u,d,s}$ and on $f_{\pi}/f_{\varphi}$. If all quark charges are set to zero $a$ is massless. For large $f_{\varphi}\gg{\rm GeV}$, $a$ becomes weakly coupled to QCD as in the usual invisible axion theories. In general, $f_{\varphi}$ is given by eq. (10). For charges $n_{q}\leq 1$, the (1,1) entry of the matrix can be rewritten as
@@ 2.3 Kaplan-Manohar ambiguity of the chiral Lagrangian
in the LO operator $\operatorname{tr}(M_{q}U^{\dagger})+\hbox{h.c.}$ [10], in view of the Cayley-Hamilton identity $2\det M_{q}\,\operatorname{tr}M_{q}^{-1}=(\operatorname{tr}M_{q})^{2}-\operatorname{tr}M_{q}^{2}$. The new term has the same symmetry transformations as $M_{q}$. The presence of $\det U=e^{i\sqrt{6}\eta^{\prime}}$ guarantees that the two terms have the same transformation also under the axial U(1)A. The constant $x_{\rm KM}$ is approximately determined by lattice computations. This is known as the Kaplan-Manohar (KM) ambiguity [10], since it complicates the extraction of bare quark masses from meson data. In particular, it replaces $m_{u}\to m_{u}+x_{\rm KM}m_{d}m_{s}$. This was considered as a loophole for reviving the massless up-quark solution to the strong CP problem, before lattice calculations established that $m_{u}\neq 0$. In our case the KM ‘ambiguity’ effectively redefines the quark mass matrix into
@@ 3 QCD-scale axion
The chiral Lagrangian at LO predicts that pions form a degenerate triplet, with the observed mass difference arising primarily from electromagnetic effects [3, 4, 5]
@@
where $\Lambda\sim m_{\rho}$. Adding a QCD-scale axion generically breaks the isospin symmetry, ruining this successful prediction. This can be seen integrating out $\varphi$ at tree level. In appendix B we present a related computation of loop corrections induced by $\varphi$ analogous to the electromagnetic splitting.
@@ 3.1 Tree-level corrections to meson masses
This class of models predicts important modifications to meson masses. To simplify formulas, we now focus on $n_{s}=0$. In the limit where the $a$ is the heaviest state, i.e. $m_{\varphi}>$ GeV, we can integrate it out. This adds to the QCD squared mass matrix of neutral mesons the following additive correction
@@ 3.2 Model with axion coupled to the up quark
Murayama [1] considered the case where the up quark only is PQ-charged, $n_{u}=1$ and $n_{d}=0$. In this case one finds (neglecting electromagnetic corrections)
@@
that is equivalent to eq. (1). A GeV-scale $\varphi$ needs a relatively large Yukawa coupling, $\kappa_{u}\gtrsim 0.5$, not much below the perturbativity limit. The KM ambiguity helps to alleviate the problem with pion masses [1], but only mildly. It shifts the quark mass matrix into
@@
The KM contribution does not modify $\langle{\varphi}\rangle$. The problematic correction to the pion mass then gets suppressed if the up entry of $M_{q}|_{\rm KM}$ is dominated by the KM shift, as it is $\varphi$-independent. Assuming that the up quark mass induced by $\langle\varphi\rangle$ is a factor $R$ smaller than the measured up quark mass, we obtain
@@
Avoiding the pion mass problem requires $R\gtrsim 10$. This corresponds to a value of $x_{\rm KM}$ much larger than the value suggested by lattice computations, that favor a mild KM shift and exclude $m_{u}=0$, corresponding to $R=\infty$ [11, 12, 2]. Thus it seems unlikely that the KM ambiguity solves the pion mass problem.
If the model were viable, the new GeV-scale scalars should be identified with some observed hadron resonances. Table 1 lists some candidates. Decay rates of $a$ and $\sigma$ can be estimated either from decays to quarks or to pions. Both computations are approximate and lead to similar estimates:
@@
The relatively narrow $a$ could be identified with the observed $\eta(1295)$, as its mass is $m_{a}\sim 1.3\,{\rm GeV}$ and its width is $\Gamma_{a}\sim 25\,{\rm MeV}$ for $\kappa_{u}\sim 1/2$ (we fixed factors of 2 compared to [1]). The broad $\sigma$ can be more easily hidden inside CP-even QCD resonances. In case the pion mass problem were eventually solved, a dedicated analysis of $a$ and $\sigma$ decays, similar to the one performed for GeV-scale axion-like particles in e.g. [13, 14, 15, 16, 17], would be needed.
@@ 3.3 Model with axion coupled to up and down quarks
Given that the pion mass splitting originates from the breaking of the SU(2) isospin symmetry, it is tempting to assume that both up and down quarks are PQ charged. For negligible $\lambda$ the mass of $\sigma$ becomes
@@
A cancellation of the pion mass shift would take place for $n_{u}=2$ and $n_{d}=1$ (see eq. (24)), as data allow for $m_{d}/m_{u}=2$ within uncertainties. However $n_{u}=2$ means that the up-quark mass arises from a dimension-5 operator that gets strongly coupled at energies just above $m_{u}$, well below the QCD scale. This would require a UV completion at that scale, with extra particles that would be problematically light.